OpenStudy (anonymous):

Solve the inequality for v 15>v -3

OpenStudy (anonymous):

Are you sure, you typed the equation correctly? [v (?)15] is it a + , -, /, or 15v?

OpenStudy (anonymous):

@ColetteB82 in order to solve I just need to know what the symbol is?

OpenStudy (anonymous):

solve for x x+4=2

OpenStudy (anonymous):

x+4=2 Subtract both sides by -4 -4 -4 x=-2

OpenStudy (anonymous):

What about you first question do you need help solving that? If so post it entirely.

OpenStudy (anonymous):

@ColetteB82 ? Are u there? Do you need more help?

OpenStudy (anonymous):

solve for w -3= 4+w

OpenStudy (anonymous):

Okay so like your first equation, your goal is to isolate the variable, and to combine like terms (3 and 4) so to move terms across the equal sign use reverse operations to solve. w-3=4+w Add 3 to both sides w=7+w Subtract w from both sides w=7 or 7

OpenStudy (anonymous):

solve for x -7= x+9

OpenStudy (anonymous):

Right lets work this one together to see if u get it! x-7=x+9 What should I do for my first step?

OpenStudy (anonymous):

subtract -7 from both sides?

OpenStudy (anonymous):

No, reverse operations. Because 7 is a negative it must be added to both sides x= x+16 x is a positive so subtract from both sides 16 The x's eliminate each other -x and a positive x (Like -1+1=0)

OpenStudy (anonymous):

Forgot to mention above, the w's cancel each other so you are left with 7 :) Do u understand it now??